Seems like there was a 5.3 magnitude EQ ~10km from us (though I didn't feel it). All watchdogs were tripped so our mirrors definitely felt it. ITMX is stuck (but all the other optics are damping fine). I tried the usual jiggling of DC bias voltage but ITMX still seems stuck. Probably a good sign that the magnet hasn't come off, but not ideal that I can't shake it free..
edit: after a bit more vigorous shaking, ITMX was freed. I had to move the bias slider by +/-10,000 cts, whereas initially I was trying +/-2000 cts. There is a tendency for the optic to get stuck again once it has been freed (while the optic's free swinging motion damps out), so I had to keep an eye out and as soon as the optic was freed, I re-engaged the damping servos to damp out the optic motion quickly.